OpenStudy (anonymous):

Choose the polynomial that is written in standard form. (4 points) 2x2y2 + 3x4y + 10x6 4x4y2 + 6x3y5 + 10x -3x8y2 + 9x2y + 10x2 -7x6y2 + x3y8 + 10x2

OpenStudy (igreen):

Standard form is writing it from highest degree to lowest degree. Do you know what a degree of a polynomial is? @ddsl

OpenStudy (anonymous):

yes but i was thinking it was B is that correct?

OpenStudy (igreen):

No, that's wrong..

OpenStudy (igreen):

\(2x^2y^2 + 3x^4y + 10x^6\) \(2x^2y^2\) degree = 4 \(3x^4y\) degree = 5 This is wrong too, because the degree of the first term is 4, and the second is 5.

OpenStudy (anonymous):

so it from highest to lowest working the prob.....................................................................

OpenStudy (igreen):

Yep^^

OpenStudy (anonymous):

i got D if its wrong i made a mistake some where

OpenStudy (igreen):

D is also wrong..

OpenStudy (igreen):

\(-7x^6y^2 + x^3y^8 + 10x^2\) \(-7x^6y^2\) degree = 8 \(x^3y^8\) degree = 11

OpenStudy (anonymous):

@#$% im not that good at these things

OpenStudy (anonymous):

now i know its C but im going to solve it

OpenStudy (igreen):

Yep.

OpenStudy (igreen):

\(-3x^8y^2 + 9x^2y + 10x^2\) \(-3x^8y^2\) degree = 10 \(9x^2y\) degree = 3 \(10x^2\) degree = 2

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @ddsl

OpenStudy (anonymous):

ok help me with one more

OpenStudy (anonymous):

Identify the Exponent Law shown in the example below. a2a3 = a2+3 = a5 Power of a Product Power of a Power Product of Powers None of the above

OpenStudy (igreen):

Sure, I'll help but you should close this question and open another one. Then tag me using the @ sign.